OpenStudy (anonymous):

Which fraction would be most likely estimated as 1/2? A 9/10 B 5/6 C 5/11 D 2/9

OpenStudy (anonymous):

I would choose C. 5/11 is closest to 6/12 which is of course 1/2.

OpenStudy (anonymous):

5/11 is closest

OpenStudy (anonymous):

no. C. bcoz 1/2=.5 9/10=.9, 5/6=.833..., 2/9=.2222...AND 5/11=.45 Which is approximately equal to .5 or 1/2. So C is the correct answer